### v1

# Building on Efficient Foundations: Effectively Training LLMs with Structured Feedforward Layers

Abstract
State-of-the-art results in large language models (LLMs) often rely on scale, which becomes computationally expensive. This has sparked a research agenda to reduce these models’ parameter count and computational costs without significantly impacting their performance. Our study focuses on transformer-based LLMs, specifically targeting the computationally intensive feedforward networks (FFN), which are less studied than attention blocks. We consider three candidate linear layer approximations in the FFN by combining efficient low-rank and block-diagonal matrices. In contrast to many previous works that examined these approximations, our study i) explores these structures from the training-from-scratch perspective, ii) scales up to 1.3B parameters, and iii) is conducted within recent Transformer-based LLMs rather than convolutional architectures. We first demonstrate they can lead to actual computational gains in various scenarios, including online decoding when using a pre-merge technique. Additionally, we propose a novel training regime, called self-guided training, aimed at improving the poor training dynamics that these approximations exhibit when used from initialization. Experiments on the large RefinedWeb dataset show that our methods are both efficient and effective for training and inference. Interestingly, these structured FFNs exhibit steeper scaling curves than the original models. Further applying self-guided training to the structured matrices with 32% FFN parameters and 2.5$\times$ speed-up enables only a 0.4 perplexity increase under the same training FLOPs. Finally, we develop the wide and structured networks surpassing the current medium-sized and large-sized Transformer in perplexity and throughput performance. Our code is available at https://github.com/CLAIRE-Labo/StructuredFFN/tree/main.
(a) Scaling curves
(b) Performance comparisons
Figure 1: (a) The scaling curves during training between the standard Transformer and the modified version with a Structured FFN, which retain 63% and 32% of the original parameters, respectively. The loss curve of the structured FFN is steeper, with a larger gap compared to the dense one at the beginning and a much smaller gap at larger sizes. (b) Performance between dense Transformer, Structured FFN with 32% parameters either trained with self-guided training or more tokens across four sizes. Circle size indicates model FLOPs. To enable straightforward comparison, we controlled their training FLOPs to be the same.
1 Introduction
Transformer language models [1] have gained significant attention for their performance and scalability. These models have grown from hundreds of millions of parameters [2] to hundreds of billions [3, 4, 5], increasing the need for efficient training and inference techniques. While much research focuses on attention, feed forward networks (FFNs) account for over 60% of the model’s parameters and FLOPs, significantly impacting latency. [^1] For example, we find that it composes 54% of total latency in a 1.3B model. Recent large-scale models [6, 7] further increase the FFN size, leading them to dominate the cost of the model compared to the attention layer.
Structured linear transformations, such as low-rank or block-diagonal matrices, are important paradigms for reducing the computational cost of feedforward layers. However, they have not yet been thoroughly explored at a sufficient scale to reduce pre-training costs and latency incurred during the inference phase in modern LLM architectures, where the main focus so far has been on improving the efficiency of the self-attention mechanism.
In this work, we investigate structured matrices for FFN blocks from the train-from-scratch aspect, first identifying their efficiency and optimization challenges and then presenting experimental results, analyzing and characterizing the behavior of models trained with structured matrices, and comparing their results. We consider three efficient linear parametrizations: LowRank, BlockShuffle (comprising two block-diagonal matrices), and BlockDense (a combination of dense and block-diagonal matrices). First, while they have demonstrated materialized computational gains, they face challenges in the practical online decoding scenario of LLM, which may process only limited input tokens at one time, leading to under-utilization of computing resources and decreased efficiency due to the additional linear projection. We address this with a pre-merge technique that restores efficiency to the original dense parametrization. Second, regarding training, we noticed that these parametrizations of the FFN blocks would suffer from either slower convergence or less stable training dynamics. To counter this, we propose a flexible and fast method we refer to as self-guided training. It employs a dense matrix as a residual component during the initial training phase, steering the training process away from suboptimal starting points gradually.
We conduct our experiments at scale on Transformers ranging from 110M to 1.3B parameters. By replacing the original FFN with efficient alternatives that use only 32% of the FFN parameters, we boost the training speed of the 1.3B model by 1.35$\times$ while only increasing the perplexity by 1 point. Interestingly, scaling curves in 1(a) indicate that structured matrices yield steeper loss scaling curves than the traditional Transformer at its optimal trade-off, suggesting a high potential for even better performance at larger scales. Furthermore, our results in 1(b) demonstrate that self-guided training enhances learning dynamics and thus the final performance, such as the 0.4 perplexity loss on the 1.3B model with the same training FLOPs and 2.5$\times$ faster FFN with 32% original parameters at inference.
Based on these findings, we design wide and structured networks at medium and large sizes using the efficient GQA [8] technique for attention and structured linear parameterizations for FFN blocks. With the same training FLOPs, our model outperforms the existing Transformer with GQA, boosting maximum throughput by 8% and 17%, respectively, while achieving slightly lower perplexity. As the first work to explore structured matrices at the scale of recent LLMs, we hope our findings and results will shed new light on the study of efficient NLP architectures.
2 Method
Multiple techniques have been proposed to approximate linear projections from sparsity to low-rank approximations. We provide an in-depth discussion of existing literature on approximating linear layers in section 3, that better contextualizes our work. We focus on structured approximations of linear projections $g({\bm{x}})={\bm{W}}{\bm{x}}$ that have the form $g({\bm{x}})={\bm{U}}({\bm{V}}{\bm{x}})$, where ${\bm{U}}$ and ${\bm{V}}$ are structured matrices, e.g. low-rank or block-diagonal. We opt for this particular form because it allows us to readily exploit the computational gains on existing hardware using existing libraries with minimal alteration. Such approximations have been previously studied in different contexts. Our contributions are exploring them (i) to approximate FFN layers of transformers, (ii) when applied from initialization, (iii) testing them at up to 1.3B scale, and providing scaling curve studies.
Figure 2: Structured linear parametrization: We show the structured linear parametrization with input dim. of $N$ and output dim. of $M$. a) The traditional dense linear parametrization. b) LowRank parametrization with a bottleneck of size $R$ where $R$ is less than $M$ and $N$. c) BlockShuffle with two block-diagonal matrices with blocks of size $B$ interleaved with a shuffle operations that mixes information from different blocks similar to ShuffleNet. d) BlockDense with the first matrix as a block-diagonal and the second a low-rank or dense matrix.
2.1 Structured linear parametrization
We explore three structures to approximate the standard linear layer ${\bm{W}}{\bm{x}}$, maintaining its input dimension $N$ and output dimension $M$ of weight ${\bm{W}}$.
LowRank
Low-rank matrices have been widely used to decompose pre-trained weights for downstream compression [9] or to construct adapters for efficient fine-tuning [10]. Researchers [11] suggest that dense layers tend to naturally converge to low-rank solutions during training, making this approximation ideal. Inspired by this, we explore low-rank matrices as alternatives to traditional linear layers, imposing this structure from initialization and investigating it during pre-training.
Formally, the low-rank approximation of a linear layer is given as ${\bm{W}}{\bm{x}}\approx{\bm{U}}^{r}({\bm{V}}^{r}{\bm{x}})$ where ${\bm{U}}^{r}\in\mathbb{R}^{M\times R}$, ${\bm{V}}^{r}\in\mathbb{R}^{R\times N}$ and $R<\min(M,N)$. Note that we use the superscript r to indicate that these matrices are used to create a low-rank approximation by projecting to or from a low-dimensional code, a notation that would become useful later on to distinguish such components from block-diagonal ones. The parameter count and MAC (Multiply-Accumulate Operations) decrease from $M\cdot N$ to $(M+N)\cdot R$.
BlockShuffle
Dao et al. [12] proposes using the Monarch decomposition of FFT, replacing the linear layer with interleaved block-diagonal and permutation matrices. An alternative motivation for such a structure can be derived from efficient convolutional designs of ShuffleNet [13] and separable convolution [14]. For simplicity, we explore the form introduced by ShuffleNet to linear layers.
The core idea of BlockShuffle is to reshape the feature dimension into two dimensions and first use a linear projection that mixes along one of these fictive dimensions, followed by a linear projection that mixes along the other. More precisely, we first reshape the input features ${\bm{x}}\in\mathbb{R}^{N}$ into $B$ blocks and apply the non-tied weight of $\frac{N}{B}\times\frac{N}{B}$ to each block, then flatten the intermediate feature. To achieve global perception, we regroup elements from different blocks into $B$ new blocks and apply the same transformation for each block again.
Technically, we can express the per-block transformation using block-diagonal matrices and formulate the above process as ${\bm{W}}{\bm{x}}\approx f^{-1}({\bm{U}}^{b}f({\bm{V}}^{b}{\bm{x}}))$, where block-diagonal matrices ${\bm{V}}^{b}$ and ${\bm{U}}^{b}$ has $B$ blocks with shapes $\frac{\min(N,M)}{B}\times\frac{N}{B}$ and $\frac{M}{B}\times\frac{\min(N,M)}{B}$ per-block. As shown in Fig. 2, the shuffle function $f(\cdot)$ enables global feature mixing by cycling different blocks and can be implemented by simply transposing and reshaping inner features. The inverse function $f^{-1}(\cdot)$ permutes the outputs back to their original order.
By separating features into two dimensions, only a few elements of the features will be processed each time. The parameter count and MAC are reduced from $M\cdot N$ to $\min(N,M)\cdot\frac{(M+N)}{B}$, where $B$ acts as a trade-off of accuracy and efficiency.
BlockDense
The previous parametrization incorporates two additional shuffle operations, which can be slow on the device. We propose an alternative efficient parametrization that retains the first block-diagonal matrix ${\bm{V}}^{b}$ but substitutes the second with a dense matrix ${\bm{U}}^{r}$. Thus, we can mix the features between blocks without permuting the inner features.
The formula is defined as ${\bm{W}}{\bm{x}}\approx{\bm{U}}^{r}({\bm{V}}^{b}{\bm{x}})$, where ${\bm{V}}^{b}$ is the block-diagonal matrix with $b$ blocks in shape $\frac{R}{B}\times\frac{N}{B}$, and ${\bm{U}}^{r}\in\mathbb{R}^{M\times R}$. Technically, the second projection does not need to be a low-rank approximation because $R$ can be larger than $M$. Nevertheless, in practice, we chose $R<M$ to limit the search space of this work, and thus use the superscript $r$ for the second matrix. The parameters of this method are determined by two variables $B$ and $R$, cutting the original burden from $M\cdot N$ to $R\cdot(M+\frac{N}{B})$. Note that BlockDense can recover the LowRank approximation if we set $B=1$ and $R<\min(M,N)$.
Implementation. We take Dao et al. [12]’s implementation for the BlockShuffle and carefully manage memory copies for BlockDense. In our experiments, we chose $B$ as a common divisor of $M$ and $N$ or $R$. Proper initialization is also investigated in Sec. B.1.
We limit our exploration of efficient linear layers within the FFN blocks.These typically have $8\cdot H^{2}$ parameters and MAC, where $H$ standards for hidden state dimension and $4\cdot H$ as the intermediate hidden size of FFN. In contrast, the proposed parametrizations have:
LowRank:  $\;10\cdot H\cdot R$ BlockShuffle:  $\;10\cdot\frac{H}{B}$ BlockDense:  $\;5\cdot H\cdot R\cdot(1+\frac{1}{B})$
2.2 Maintaining efficiency during online decoding
Parallelism-bound FFN With reduced FLOPs and parameters, our proposed linear parametrization can accelerate the model for compute-bound and memory-bound scenarios [15], usually during training, prefilling, and decoding with a relatively big batch size. However, for online decoding with a very small batch size and sequence length of 1, a practical scenario for LLM, both FFN and structured FFN can become parallelism-bound [16] with poor utilization of the GPU resources, especially on powerful devices like A100. Because each linear projection suffers from parallelism-bound, efficient linear parametrization may lead to worse latency performance due to doubling the number of linear projections. We propose a pre-merge technique to mitigate the issue.
Pre-merge technique Taking advantage of the fact that these parametrizations do not have non-linearity, we propose to combine the structured matrices into a single dense layer and keep both the structured and the dense one for online decoding. Then, we can dynamically decide which parametrization to use based on the current batch size and setting. Fig. 5 analyzes using structured or dense forms for different batch and model sizes, allowing us to decide when to use the pre-merged linear projection.
2.3 Addressing the optimization challenge
Using the efficient parametrization from initialization can suffer from optimization difficulty because the deep linear parametrization introduces additional symmetries [^2] Symmetries arise because such a factorization ${\bm{U}}{\bm{V}}$ is not unique; for any invertible matrix ${\bm{C}}$ of size $R\times R$ we have ${\bm{U}}{\bm{V}}={\bm{U}}{\bm{C}}{\bm{C}}^{-1}{\bm{V}}=({\bm{U}}{\bm{C}})({\bm{C}}^{-1}{\bm{V}})=\tilde{{\bm{U}}}\tilde{{\bm{V}}}$. BlockShuffle can also be included by fusing the shuffle operation into block-diagonal matrices. , which is a source of proliferation of saddle points and generally less smooth loss function as pointed out in [17, 18]. We hypothesize that this makes poorer learning dynamics of the structured parametrization. Empirically, we show that the deep linear form of ${\bm{U}}({\bm{V}}{\bm{x}})$ in Fig. 3 leads to instability and loss spike or to slow convergence compared to the dense linear projection. We further elaborate on this, highlighting how the inconsistency of gradient updates between the structured parametrization and original linear projection affect learning dynamics in Appendix A.2.
(a) Loss spikes
(b) Slower convergence
Figure 3: Poor training dynamics: Training dynamics of LowRank with rank of 128 under different training configurations. Curves correspond to a 4-layer Transformer with a model width of 768 on WikiText-103. We apply self-guided training in the first half of training. Refer to Sec. A.1 for additional details.
Self-guided training Addressing the poor training dynamics by carefully tuning the learning rate schedule and gradient clipping coefficient might be possible, but it is costly and may switch between slow convergence and training instability. We propose a less costly and simple approach that can be used with minimal re-tuning of hyperparameters.
To motivate our proposal, we start by finding that the updates on ${\bm{U}}{\bm{V}}$ scales the function of the backpropagated gradients ${\bm{g}}$ (see App. A.2), then turn into the typical training dynamics with gradients. An issue that needs to be addressed in the early stage of training is feature specialization when the learning process assigns semantics to the different hidden units of the model, sometimes also phrased as identifying the winning ticket [19]. In this process, certain weights will need to be suppressed, and symmetries in the parametrization must be resolved.
To address this problem, we propose using the dense parametrization ${\bm{W}}$ as a crutch to efficiently make decisions about feature specialization and then transfer them to ${\bm{U}}$ and ${\bm{V}}$ through ${\bm{g}}$. To this end, we use the following parametrization
${\bm{o}}=\alpha\cdot{\bm{W}}{\bm{x}}+(1-\alpha)\cdot{\bm{U}}({\bm{V}}{\bm{x}}).$ (1)
${\bm{o}}$ is the layer’s output, and $\alpha$ decays following a cosine scheduler. As a residual component, learning ${\bm{W}}$ is unaffected by the additional saddles and pathologies introduced by the structured parametrization, allowing units to specialize. This guides the training of ${\bm{U}}$ and ${\bm{V}}$, which are forced slowly to take over by providing the hidden units semantics learned by ${\bm{W}}$. This approach also relates to homotopy methods such as simulated annealing, where a hard optimization problem is transformed into an easier to optimize form with certain desired properties, gradually transforming the problem to its original form. Here, we consider the easier optimization problem is to train with dense matrices. By decreasing alpha from $1$ to 00, we transform the easier-to-optimize loss into the original parametrization we want to use.
Furthermore, we initialize ${\bm{W}}_{0}={\bm{U}}_{0}{\bm{V}}_{0}$, making the observation that by using this initialization for the dense residual branch, we can easily start the guided training at any stage of learning (e.g., fine-tuning) without affecting the model behavior. We refer to this as self-guiding training, as the better learning dynamics from ${\bm{W}}$, which initially is just ${\bm{U}}{\bm{V}}$, guide the learning of ${\bm{U}}$ and ${\bm{V}}$ through the backpropagated gradients ${\bm{g}}$.
Reducing the computational cost of self-guided training: Note that while $\alpha>0$, we need to perform forward and backward passes through the dense version of the weight matrix ${\bm{W}}$, which could be expensive. To address this issue, we consider the following stochastic version of the above formula, which allows us to control how often we need to use the dense residual branch:
${\bm{o}}=\begin{cases}\alpha\cdot{\bm{W}}{\bm{x}}+(1-\alpha)\cdot{\bm{U}}({\bm{V}}{\bm{x}}),&p<\alpha\\ {\bm{U}}({\bm{V}}{\bm{x}}),&p\geq\alpha.\end{cases}$ (2)
In our practice, $p$ is a random variable sampled independently from a uniform distribution over 0 to 1 in each training forward pass. With $\alpha$ following a cosine annealing schedule, this softer version Eq. (missing) 2 reduces the expected computation to half of Eq. (missing) 1. For example, using our method for half the training time increases the FLOPs by only 25% of the original FFN. This has a negligible impact on accuracy. More ablation studies of this technique are presented in Sec. B.2.
3 Related work
Efficient techniques for training LLMs Recent advancements in attention mechanisms have significant improvements in the efficiency of attention [20, 21, 22, 8, 23, 24, 25], and the focus has shifted towards improving FFNs in LLMs, which contributes to at least half of the training time. Dynamic architectures such as mixtures of experts [26, 27, 28], or optimizers with faster convergence [29, 30] have been popular in improving training efficiency. Moreover, Xia et al. [31] employs structured pruning with learned sparsity masks and a dedicated data-loading policy to reduce the training budget.
There has been a recent focus on parameter-efficient fine-tuning methods like LoRA [10] and structured approximation of the linear layers (see, [9]). LoRA uses the low-rank approximation to reduce trainable parameters during the finetuning phase, whereas Sharma et al. [9] selectively applies low-rank decomposition to well-trained weights. While these methods used low-rank approximation of the weights, they did not focus on pre-training.
Structured matrices in deep learning Researchers use structured matrices in the form of dense matrices with shared parameters, like Circulant and Toeplitz [32] [^3] We ran preliminary experiments using Circulant, Toeplitz matrices, and convolutions to improve the efficiency of FFNs, but our initial results were negative (slow and worse performance) and we did not pursue this direction further., and structured matrices, such as low-rank and diagonal, to reduce parameters and FLOPs while optimizing CUDA kernel use. Low-rank matrices, initially used in convolutional networks [33], have shown high efficiency in training [34], achieving up to a 2.9$\times$ speed-up with similar performance. Some studies [35, 36] adapt the rank during training and suggest regularizers to maintain SVD decomposition for better accuracy. Khodak et al. [37] propose spectral initialization and aligning weight decay of matrix products with standard linear layers. However, these studies mainly focus on ResNets [38] rather than recent LLMs. There have been other studies that aim to improve the expressiveness of structured matrices. For instance, Moczulski et al. [39] uses interleaved diagonal and Fourier transforms, while Dao et al. [40] proposes butterfly parametrization for various transformations. These approaches often lack efficiency due to additional layers or irregular sparsity patterns. Dao et al. [12] simplified butterfly matrices to block-diagonal ones, achieving a 2$\times$ speed-up on WikiText-103 language modeling tasks. In this work, for accuracy and efficiency, we explored low-rank factorization of weight matrices with reduced bottleneck dimension and block-diagonal matrices to reduce parameters in our LLM training studies.
4 Experiments
In our experiments, we analyzed the scaling curves, efficiency, and training challenges of structured parametrization of FFN module in LLMs.
Model We use a basic Transformer architecture [41, 2] with Rotary Embedding [42] and SentencePiece Tokenizer [43]. The original FFN module includes two linear layers and a GeLU activation. Our model sizes include 110M (Transformer-s), 335M (Transformer-m), 729M (Transformer-l) and 1.3B (Transformer-xl) parameters, with sequence lengths of 1024 and model widths from 768 to 2048. We only replace FFN modules with efficient linear layers, as the attention module is well-studied [8, 22]. We explore two alternatives to dense FFNs, reducing parameters to 63% or 32% of the original size by adjusting rank and block number. The first FFN module is not replaced to avoid non-negligible performance degradation.
Training We use the RefinedWeb dataset [44] and randomly select 0.5B tokens for validation, reserving the rest for training. Following scaling laws [45], we allocate tokens at 20 times the number of each baseline model’s parameters. We train using A100 80G GPUs with mixed precision (bfloat16 and float32) and adopt fast CUDA kernels like Flash Attention [24] in all experiments for fair comparisons. Hyperparameters, including learning rates and the global number of tokens per batch, are set according to recent studies [46, 47]. We measure training FLOPs as in Megatron [48], including all matrix multiplications. For more details on our experimental settings, see Appendix C.
Figure 4: Latency and FFN sizes: Latency over 30K tokens of different FFN sizes. The intermediate size of the FFN is set to be 4 times the FFN width. In (a), we show latency results of 63% and 32% FFN parameters across different FFN widths. (b) shows the latency of LowRank, BlockShuffle, and BlockDense with different numbers of parameters by varying the rank and number of blocks from the dense FFN with a width of 5120.
Figure 5: Latency over different batch size for different widths: Decoding latency results between dense FFN and structured matrices with 32% FFN parameters across different widths and batch sizes. Note that we have a sequence length of $1$ at the decoding phase; thus, $T$ equals batch size.
Table 1: Performance comparisons of different FFN parameterizations: Performance of two sizes of different structured matrices with 63% and 32% of the original FFN module’s parameters. We report model size, total FFN size, training tokens, training FLOPs, and training time. Note that the total structured FFN is not exactly 63% of the original because we don’t replace the first FFN module. Also, the BlockDense is slightly smaller for -m and -xl models to ensure the rank is a multiple of 256 when matching parameters. Loss and perplexity (PPL) are evaluated on a 0.5B token validation set. Please refer to Table 9 for the whole table.
Architecture Model FFN Training PPL Size (M) Size (M) Tokens (B) FLOPs Time (h) Transformer-xl 1274 805 25.5 2.10e+20 352.2 12.46 63% LowRank (R=1024) 985 516 25.5 1.66e+20 302.2 12.86 BlockDense (B=4, R=1536) 955 486 25.5 1.61e+20 298.7 12.97 BlockShuffle (B=2) 985 516 25.5 1.66e+20 330.6 12.98 32% LowRank (R=512) 744 275 25.5 1.29e+20 260.2 13.55 BlockDense (B=4, R=768) 728 260 25.5 1.27e+20 261.2 13.74 BlockShuffle (B=4) 744 275 25.5 1.29e+20 284.9 13.81
4.1 Evaluating latency results
We investigate the efficiency of structured FFN and consider different numbers of tokens $T$ to discuss different scenarios. Here, $T$ corresponds to the total number of tokens in a batch.
Large number of tokens Using large $T$, the standard linear layers and our efficient structured parametrizations become computation-bound where FLOPs become a latency bottleneck [15]. This setting mainly concerns training, the prefill phase of inference, and extensive offline decoding. By reducing FLOPs, it can be seen that in 4(a) and 4(b), our designs can improve the latency consistently. In 4(b), we showed the acceleration of structured parametrization for an FFN module across different ranks or blocks. For BlockDense, the block number does not affect latency performance, all achieving a 1.72x speed-up with 50% fewer parameters. 4(a) compares FFN and different structured FFN results with 63% and 32% parameters across different widths. GPU utilization improves with increasing width. From width 1536, LowRank and BlockDense with 63% FFN parameters achieve a 1.4$\times$ speed-up and a 2.5$\times$ speed-up. BlockShuffle offers modest improvements, with 1.1$\times$ and 2.0$\times$ speed-ups for the two cases, respectively.
Small number of tokens FFN can be parallelism-bound with small $T$ on the A100 GPUs, for example, when $T=16$. With increasing $T$, FFN becomes memory-bound and will eventually be computation-bound. Small $T$ can occur during online and offline decoding stages when unrolling the model step by step. As discussed earlier, our pre-merge method can alleviate the parallelism-bound issue and maintain the same latency with dense matrices. Fig. 5 shows the latency results for three different widths, varying the batch of tokens to determine when to use efficient alternatives or choose pre-merged dense matrices. For example, with a 2048-width FFN, it is difficult to fully utilize resources on GPU with limited tokens. The performance improves significantly when using width 5120 and 6144, such as speed improvements of 2.63$\times$ speed-up of LowRank with 32% FFN parameters on $T=2048$ and 2.81$\times$ acceleration of BlockDense with 32% parameters on $T=1536$.
4.2 Findings on efficient training
(a) BlockDense
(b) BlockShuffle
Figure 6: Scaling curves of structured matrices: The scaling curves during training between the standard Transformer and the modified version with a Structured FFN, which retains 63% and 32% of the original parameters, respectively.
Here, we analyze the performance of Transformers using FFN with 63% or 32% of the original FFN parameters by reducing the matrix rank and number of blocks. Experiments are conducted with various Transformer sizes, comparing performance among different structured matrices and showing the scaling curves.
Comparison between structured FFNs We conducted experiments fixing the model and training FLOPs for LowRank and BlockDense when comparing them to BlockShuffle in Sec. 4 and Table 9. We achieved a 0.8 perplexity reduction on Transformer-s and a 0.4 perplexity reduction on Transformer-m. We think that for FFNs in language models, BlockShuffle may not be an ideal option. However, in vision tasks where ShuffleNet has demonstrated success and block-diagonal matrices are better for local information, we show in Sec. B.3 that the block-diagonal matrix is a more suitable inductive bias. Moreover, LowRank and BlockDense, with 63% of the parameters in FFN, increase the loss by only 0.02 to 0.04 while reducing the total training time by about 15% on Transformer-xl and Transformer-l. Additionally, they can accelerate training by 1.36 times with only a 1 perplexity increase on Transformer-xl with 32% of the parameters.
Scaling analysis We compare the performance of different model sizes and show their scaling curves in 1(a), 6(a), 6(b). All the efficient parametrization gets closer to the baseline with the increasing model size. For example, LowRank with 32% parameters has a $0.08$ loss gap to Transformer-xl compared to a loss gap of $0.12$ at the scale of Transformer-s models.
According to 1(a), we observe that: (i) The structured matrices exhibit steeper scaling curves compared to the dense networks, indicating significant potential for these efficient designs in LLMs. (ii) The scaling curve of 32% parameters of FFN is steeper than the 63% parameters of FFN highlights the scaling potential of highly structured large models. (iii) Given fixed training FLOPs budget, a wider and structured network with more tokens may achieve comparable or superior performance to dense networks at the optimal trade-off.
Our scaling curves for the structured matrices can be further optimized: (1) they are not drawn at their optimal training-compute trade-off while the baseline is. (2) Only the FFN is made structured, while attention remains dense and thus contributes more to the whole model, which is less optimal. The second point also explains why the current 32% parameter curve shows a larger validation loss than the 63% parameter curve and motivates us to reduce attention further using existing techniques in Sec. 4.4.
4.3 Self-guided training
We apply self-guided training during the first half of training to demonstrate its effectiveness. As shown in Sec. 4.3 and Sec. B.2, our method consistently reduces loss across all efficient parametrizations, improving the perplexity by 1.2 for Transformer-s and 0.8 for Transformer-m. Then, to enable a straightforward comparison under the same training FLOPs, we adjust the training steps for self-guided training and repeat those tokens at the end to ensure they’re fully learned by structured matrices. As can be seen in Sec. B.5 and 12(a), 1(b), 12(b), this reduces the perplexity gap for Transformer-xl from 1.0, 1.2, and 1.3 to 0.5, 0.5, and 0.6 for LowRank, BlockDense, and BlockShuffle, respectively, under the same training FLOPs and 32% inference FFN FLOPs. Additionally, we compare our method with another advanced baseline that trains structured parametrizations with more tokens, showing that ours achieves comparable or superior results even with the same number of tokens.
Table 2: Wide and structured network performance: For GQA, we increase the intermediate FFN size as in [6] to match parameters. For ours, we either use more tokens or self-guided training (marked with ) to match the training FLOPs. TP refers to the maximum throughput measured on a generation length of 256. Detailed configuration is put in Table 12.
[…]
Table 3: Ablation study about self-guided training across methods. We apply it to the first half of training with 25% increased FFN budget. For more results, please refer to Sec. B.2.
Architecture FFN Training FLOPs PPL Transformer-m 201M 1.55e+19 18.29 LowRank 69M 1.01e+19 20.60 LowRank 1.21e+19 19.90 BlockDense 65M 1.00e+19 20.85 BlockDense 1.19e+19 20.10 BlockShuffle 69M 1.01e+19 21.12 BlockShuffle 1.21e+19 20.36
4.4 Wide and structured network
In the previous section, we explored structuring FFN, achieving steeper scaling curves and better results via self-guided training. As maintaining the parameter ratio of attention to FFN can be important, in this section, we use GQA [8] to make attention efficient and LowRank for FFN, designing a wide and structured network from Transformer-m and Transformer-l (see Table 2). To match the training FLOPs, we either train on more tokens or apply self-guided training. This approach aims to enhance efficiency with a much smaller network, achieving an 8% and 17% maximum throughput boost on Transformer-m and Transformer-l, respectively, while maintaining or slightly improving perplexity.
5 Conclusion
In this paper, we conducted extensive experiments investigating the use of structured matrices to parameterize FFN in Transformers, with models up to 1.3B parameters on the RefinedWeb dataset. Our primary aim was not to determine which structured matrices perform best, as this can be task-dependent, but to explore common issues including efficiency and optimization challenges of existing structured matrices as well as BlockDense. For example, with the proposed self-guided training, our models can achieve a 1.36× inference speed-up with only a 0.4 perplexity increase. With wide and structured networks, we show that we can improve both throughput and perplexity.
Limitations: BlockDense and BlockShuffle are more complicated than LowRank. In this work, we only explored a limited range of hyperparameter settings of them. However, since these approaches are new, we believe that further performance improvements may be possible by better tuning their hyperparameters. We primarily focused on language modeling with limited vision experiments included in the appendix. Additionally, we did not explore the optimal scaling laws for structured matrices, which may further enhance performance. We hope this work will encourage further exploration of other structured linear parameterizations.
Acknowledgments and Disclosure of Funding
We are grateful to Soham De for the insightful discussions and his valuable feedback on our work.
Appendix
Appendix A Supplementary of optimization challenges
A.1 Observation of training dynamics
We provide additional training dynamics curves, including BlockDense, BlockShuffle in Fig. 8 and Fig. 9, highlighting their loss spikes or slower convergence. In contrast, we show the improved loss curves and dynamics from applying self-guided training to the worst case.
(a) $lr=1.5e-3$
(b) $lr=1.0e-3$
(c) $lr=5.0e-4$
(d) $lr=1.5e-3$
(e) $lr=1.0e-3$
(f) $lr=5.0e-4$
Figure 7: (a-c): Training dynamics of LowRank with different ranks and the dense model under different hyper-parameters. Data points are measured on a 4-layer Transformer with model width 768 and WikiText-103. We zoom into the beginning of training for clearer observations. (d-f): Training dynamics of LowRank and the self-guided training. The self-guided training overcomes the loss spikes and makes the training faster. We show the whole training curve to indicate its success. $R$ indicates the rank of low-rank matrices.
(a) $lr=1.5e-3$
(b) $lr=1.0e-3$
(c) $lr=5e-4$
(d) $lr=1.5e-3$
(e) $lr=1.0e-3$
(f) $lr=5e-4$
Figure 8: (a-c): Training dynamics of BlockDense with 2 blocks and different ranks and the dense model under different hyper-parameters. (d-f): For BlockDense ($B=2,R=128$), training dynamics of self-guided training indicated by . Other settings follow Fig. 7.
(a) $lr=1.5e-3$
(b) $lr=1.0e-3$
(c) $lr=5e-4$
(d) $lr=1.5e-3$
(e) $lr=1.0e-3$
(f) $lr=5e-4$
Figure 9: (a-c): Training dynamics of BlockShuffle with different numbers of blocks and the dense model under various hyper-parameters. (d-f): For BlockShuffle ($B=8$), loss curves of self-guided training indicated by . Other settings follow Fig. 7. Note that the training dynamics of different structured matrices are not comparable here because their sizes are not controlled to be the same.
Meanwhile, we report the loss spikes found on the Transformer-xl and RefinedWeb dataset in Fig. 10.
(a) LowRank on Transformer-xl
(b) BlockDense on Transformer-xl
Figure 10: Loss curves of Transformer-xl. Structured FFN with 32% parameters exhibits slower convergence. Also, there exist loss spikes.
A.2 On the poorer training dynamics
We observed that loss spikes occur along with large gradient norms. This motivates us to analyze the gradient updates during the backward pass of the linear ${\bm{U}}{\bm{V}}$. Considering ${\bm{g}}$ as the gradient of output and ${\bm{x}}$ as the input, the standard linear layer ${\bm{W}}$ gradient update is ${\bm{g}}{\bm{x}}^{\top}$. For the structured parametrization, the gradients of ${\bm{U}}$ and ${\bm{V}}$ are ${\bm{g}}{\bm{x}}^{\top}{\bm{V}}^{\top}$ and ${\bm{U}}^{\top}{\bm{g}}{\bm{x}}^{\top}$, respectively. Then, for ${\bm{W}}^{\prime}$ as being the updated parameters, we will have the updates for ${\bm{W}}$ to be:
$\Delta{\bm{W}}={\bm{W}}^{\prime}-{\bm{W}}=-lr\cdot{\bm{g}}{\bm{x}}^{\top}.$ (3)
And the one for ${\bm{U}}{\bm{V}}$:
$\Delta({\bm{U}}{\bm{V}})=-lr\cdot({\bm{U}}{\bm{U}}^{\top}{\bm{g}}{\bm{x}}^{\top}+{\bm{g}}{\bm{x}}^{\top}{\bm{V}}^{\top}{\bm{V}})+O(lr^{2})$ (4)
Thus, it can be seen from Eq. (missing) 4 that the projections ${\bm{U}}{\bm{U}}^{\top}$ and ${\bm{V}}^{\top}{\bm{V}}$ can disrupt the gradient ${\bm{g}}{\bm{x}}^{\top}$. If the norm of ${\bm{U}}$ and ${\bm{V}}$ are small, the new update vanishes faster than the original update, and in reverse, if their norm is large, the update blows up, leading to unstable training.
To be specific, we calculate the spectral norm of ${\bm{U}}{\bm{U}}^{\top}$ and ${\bm{V}}^{\top}{\bm{V}}$ and use this to indicate the maximum scale the matrix can stretch a vector. Fig. 11 shows that the largest singular value can vary significantly, being either much greater than or less than 1, depending on the weight shape (input dimension, rank, number of blocks) and magnitude. Interestingly, very structured FFNs with small ranks or many blocks tend to have smaller spectral norms, while others have larger ones. This corresponds to the phenomenon in Fig. 8 and Fig. 9, where smaller FFN are prone to slower convergence and larger ones to loss spikes.
An alternative intuitive perspective of this issue is to realize that learning with structured matrices has additional redundant degrees of freedom brought by the symmetry. For example, to increase the norm of a feature, it can increase either the corresponding weights ${\bm{U}}$ or ${\bm{V}}$. Given that gradient descent makes this decision independently, it will overshoot and make learning less well-behaved.
(a) N=768, std=0.02
(b) N=1536, std=0.02
(c) N=1536, std=0.05
(d) N=1536, std=0.05
(e) N=768, std=0.02
(f) N=1536, std=0.02
(g) N=1536, std=0.05
(h) N=1536, std=0.05
Figure 11: Spectral norm of the matrix ${\bm{V}}^{\top}{\bm{V}}$, where in (a-c), ${\bm{V}}$ is the low-rank matrix and in (d-f), ${\bm{V}}$ is the block-diagonal matrix. N presents the input dimension of the weight. Std indicates the standard deviation value of the normal distribution from which we sample weight elements.
Appendix B Supplementary experiment
B.1 Design choice
For the BlockDense, we also investigate the opposite order that places the low-rank or dense matrix first, followed by the block-diagonal matrix. However, changing the order does not result in good performance. For example, it increases the perplexity on the RefinedWeb dataset from 29.17 to 29.65 on Transformer-s.
For initialization, we follow the spectral initialization for LowRank, as suggested by prior works [37]. For BlockDense and BlockShuffle, motivated by Eq. (missing) 4, we propose using orthonormal initialization, setting the singular values of ${\bm{U}}_{t}{\bm{U}}_{t}^{\top}$ and ${\bm{V}}_{t}{\bm{V}}_{t}^{\top}$ to 1 at the start. Experimentally, this stabilizes training dynamics and improves the perplexity performance (Table 5). For weight decay, we tried the Frobenius decay proposed by Khodak et al. [37]; however, it did not have a clear benefit in our experiments and increased training FLOPs slightly. Hence, we adopted standard weight decay for all the structured FFNs.
B.2 Ablation of self-guided training
Table 4: Ablation study of self-guided training indicated with for Structured FFN with 32% parameters. Model FLOPs are calculated on one sample with 1024 sequence length.
Architecture Model FFN Model Training Loss PPL Size (M) Size (M) FLOPs (G) Tokens (B) Steps (K) FLOPs Transformer-s 110 56.62 262.9 2.2 4K 1.69e+18 3.2569 25.97 LowRank 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3748 29.22 LowRank 2.2 4K 1.39e+18 3.3329 28.02 BlockDense 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3731 29.17 BlockDense 2.2 4K 1.39e+18 3.3338 28.04 BlockShuffle 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3994 29.95 BlockShuffle 2.2 4K 1.39e+18 3.3583 28.74 Transformer-m 335.08 201.33 788.7 6.7 13K 1.55e+19 2.9062 18.29 LowRank 202.43 68.68 517.0 6.7 13K 1.01e+19 3.0251 20.60 LowRank 6.7 13K 1.21e+19 2.9907 19.90 BlockDense 198.67 64.91 509.3 6.7 13K 1.00e+19 3.0371 20.85 BlockDense 6.7 13K 1.19e+19 3.0008 20.10 BlockShuffle 202.43 68.68 517.0 6.7 13K 1.01e+19 3.0501 21.12 BlockShuffle 6.7 13K 1.21e+19 3.0135 20.36
Experiments in this part are conducted on Transformer-s and Transformer-m with ranks 192 and 256 in Table 6, respectively. We apply self-guided training during the first half of the training process.
First, we compare stochastic self-guided training with the static version. The stochastic and faster version in both model sizes brings about a 0.1 perplexity increase while reducing computation by half. Second, other techniques are compared. Dense layer decomposition, which decomposes the weight directly at the midpoint of training, is examined. This approach can lead to abrupt loss increases in training curves, resulting in worse performance. Strategies incrementally reducing rank require a feasible and complex change strategy and fail to address the inconsistent gradient update problem, thus still suffering from poor results Table 6.
Generally speaking, our method stands out due to its flexibility, simplicity, and efficiency. Eq. (missing) 1 makes it adaptable to any efficient linear parametrization without special constraints, while progressive rank reduction and direct decomposition require a feasible solution to evolve. Our guided initialization allows its usage in various stages of training without the need for a well-trained teacher. It is simple because Eq. (missing) 1 provides only one smooth transition. It is efficient due to the layer-specific definition in Eq. (missing) 1 and stochastic computation in Eq. (missing) 2.
Table 5: Different initialization of BlockShuffle and BlockDense, where random indicates random Gaussian initialization and orthonormal indicates orthonormal initialization. Data points are measured on the 4-layer Transformer and WikiText-103 with a learning rate of 1.0e-3.
Method Initialization PPL Transformer (4-layers) random 23.24 BlockShuffle (B=4) random 27.24 BlockShuffle (B=4) orthonormal 25.33 BlockDense (B=2, R=128) random 28.25 BlockDense (B=2, R=128) orthonormal 26.63
Table 6: Ablation study of self-guided training on LowRank trained on RefinedWeb.
Method PPL Transformer-s 25.97 Direct decomposition 28.56 Progressive decreasing rank 29.35 Self-guided training 28.02 Self-guided training (slower) 27.90 Transformer-m 18.29 Self-guided training 19.90 Self-guided training (slower) 19.81
B.3 Vision tasks
Table 7: Experiments on CIFAR10 and vision models, where the locality is highly preferred. The first layer column in the table indicates whether to apply structured matrices to the first FFN.
Method Structured first FFN Dense first FFN #Params (M) Accuracy #Params (M) Accuracy 5-layer MLP (H=768) 4.14 66.99 4.14 66.99 LowRank (R=192) 1.63 64.04 3.26 65.42 BlockShuffle (B=4) 1.63 67.08 3.26 65.67 ViT (H=384) 21.34 92.49 21.34 92.49 LowRank (R=24) 8.29 89.56 9.38 92.09 BlockShuffle (B=16) 8.29 90.42 9.38 92.49
Although in our main experiments, the BlockShuffle performs the worst with two block-diagonal matrices, we provide experiments on the CIFAR10 dataset here, showing that when locality is highly preferred, block-diagonal matrices may perform better than low-rank matrices.
In Table 7, experiments are conducted on 5-layer MLP (MultiLayer Perceptron) and ViT models. The 5-layer MLP consists of a linear layer, batch normalization [49], and ReLU activation function with a hidden dimension of 384. It is trained for 500 epochs with a learning rate of 1.0e-3 and a batch size of 128. For the ViT models, 12 layers with a hidden dimension of 384 are used, and they are trained for 750 epochs with a learning rate of 6.0e-4 and a batch size of 512.
Since the first layer in vision tasks typically prefers from the locality, especially in a 5-layer MLP where the image pixels are directly flattened into the input, we conducted experiments with and without structuring the first layer as LowRank and BlockShuffle. Both setups, particularly when structuring the first linear layer, show that block-diagonal matrices can be advantageous for vision tasks. Specifically, using a block-diagonal matrix to replace the first layer of the 5-layer MLP model often results in superior performance because the block structure groups neighboring pixels together, compensating for the MLP’s lack of locality. However, structuring the first FFN in ViT can lead to significant accuracy degradation, supporting our decision not to apply structured FFNs to the first layer in the main paper.
B.4 Downstream tasks
We also use the Transformer-xl trained on 25.5B tokens for downstream tasks, including PIQA, HellaSwag, Winogrande, and ARC tasks. The evaluation code is based on lm-evaluation-harness [^4] https://github.com/EleutherAI/lm-evaluation-harness with the default prompt. Note that since we train the largest model on a limited number of tokens compared to the 300B tokens used in recent papers [46, 50], the models may not perform optimally on downstream tasks in a zero-shot setting. However, our aim here is to compare the downstream performance of structured FFN and standard FFN.
From Sec. B.4, it can be seen that with structured FFN using 63% of the parameters, we observe an average accuracy degradation of 0.6 to 0.9 on these downstream tasks. With 32% of the parameters, the accuracy performance decreases by about 1.2 to 1.7 percent. While our model is trained on a limited number of tokens and evaluated with default prompts, we conjecture that training on a large enough dataset or finding prompts more suitable for structured FFN might even boost the performance.
Table 8: Performance of Transformer-xl on downstream tasks under the zero-shot setting, where PPL refers to the validation PPL on RefinedWeb dataset.
Architecture Model Size FFN Size PPL PIQA HellaSwag Winogrande ARC (easy) Avg. Transformer-xl 1274M 805M 12.46 70.46 36.62 51.62 46.30 51.25 63% LowRank (R=1024) 985M 516M 12.86 70.51 35.66 50.36 46.21 50.69 BlockDense (B=4, R=1536) 955M 486M 12.97 69.48 36.01 51.46 45.24 50.55 BlockShuffle (B=2) 985M 516M 12.98 68.61 35.54 50.99 46.34 50.37 32% LowRank (R=512) 744M 275M 13.55 68.66 34.73 50.59 44.23 49.55 BlockDense (B=4, R=768) 728M 260M 13.74 69.04 34.37 52.64 44.15 50.05 BlockShuffle (B=4) 744M 275M 13.81 68.61 33.94 50.75 44.23 49.38
B.5 Complete results table
Table 9: Performance of two sizes of different structured matrices with 63% and 32% of the original FFN module’s parameters. We report model size, total FFN size, training tokens, training FLOPs, and training time. Note that the total structured FFN is not exactly 63% of the original because we don’t replace the first FFN module. Also, the BlockDense is slightly smaller for -m and -xl models to ensure the rank is a multiple of 256 when matching parameters. Loss and perplexity are evaluated on a 0.5B token validation set.
Architecture Model FFN Training Loss PPL Size (M) Size (M) Tokens (B) FLOPs Time (h) Transformer-s 110 56.62 2.2 1.69e+18 4.0 3.2569 25.97 63% LowRank (R384) 90.17 37.16 2.2 1.44e+18 3.8 3.3017 27.16 BlockDense (B2R512) 90.17 37.16 2.2 1.44e+18 3.8 3.3034 27.20 BlockShuffle (B2) 90.17 37.16 2.2 1.44e+18 4.2 3.3191 27.63 32% LowRank (R192) 73.95 20.94 2.2 1.22e+18 3.6 3.3748 29.22 BlockDense (B2R256) 73.95 20.94 2.2 1.22e+18 3.5 3.3731 29.17 BlockShuffle (B4) 73.95 20.94 2.2 1.22e+18 4.0 3.3994 29.95 Transformer-m 335.08 201.33 6.7 1.55e+19 32.5 2.9062 18.29 63% LowRank (R512) 262.73 128.97 6.7 1.26e+19 29.6 2.9508 19.12 BlockDense (B4R768) 255.19 121.44 6.7 1.23e+19 29.9 2.9581 19.26 BlockShuffle (B2) 262.73 128.97 6.7 1.26e+19 33.1 2.9622 19.34 32% LowRank (R256) 202.43 68.68 6.7 1.01e+19 26.9 3.0251 20.60 BlockDense (B4R384) 198.67 64.91 6.7 1.00e+19 27.1 3.0371 20.85 BlockShuffle (B4) 202.43 68.68 6.7 1.01e+19 30.0 3.0501 21.12 Transformer-l 729.11 452.98 14.6 7.03e+19 130.5 2.6594 14.29 63% LowRank (R768) 566.32 290.19 14.6 5.61e+19 113.6 2.6957 14.82 BlockDense (B2R1024) 566.32 290.19 14.6 5.61e+19 114.3 2.7038 14.94 BlockShuffle (B2) 566.32 290.19 14.6 5.61e+19 124.3 2.7021 14.91 32% LowRank (R384) 430.66 154.53 14.6 4.42e+19 100 2.7527 15.69 BlockDense (B2R512) 430.66 154.53 14.6 4.42e+19 100.9 2.7570 15.75 BlockShuffle (B4) 430.66 154.53 14.6 4.42e+19 110.3 2.7735 16.01 Transformer-xl 1274.14 805.31 25.5 2.10e+20 352.2 2.5226 12.46 63% LowRank (R1024) 984.73 515.90 25.5 1.66e+20 302.2 2.5541 12.86 BlockDense (B4R1536) 954.59 485.75 25.5 1.61e+20 298.7 2.5628 12.97 BlockShuffle (B2) 984.73 515.90 25.5 1.66e+20 330.6 2.5633 12.98 32% LowRank (R512) 743.56 274.73 25.5 1.29e+20 260.2 2.6062 13.55 BlockDense (B4R768) 728.49 259.65 25.5 1.27e+20 261.2 2.6204 13.74 BlockShuffle (B4) 743.56 274.73 25.5 1.29e+20 284.9 2.6254 13.81
Table 10: Performance of self-guided training indicated by $\clubsuit$ on Structured FFN with 32% parameters under the same training FLOPs. We also include the structured FFN trained on more different tokens as a highly advanced baseline. Model FLOPs are calculated on one sample with 1024 sequence length.
Architecture Model FFN Model Training Loss PPL Size (M) Size (M) FLOPs (G) Tokens (B) Steps (K) FLOPs Transformer-s 110 56.62 262.9 2.2 4K 1.69e+18 3.2569 25.97 LowRank 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3748 29.22 LowRank 3.0 6K 1.69e+18 3.2928 26.92 LowRank 2.2 5K 1.69e+18 3.2866 26.75 BlockDense 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3731 29.17 BlockDense 3.0 6K 1.69e+18 3.2982 27.06 BlockDense 2.2 5K 1.69e+18 3.2856 26.73 BlockShuffle 73.95 20.94 189.8 2.2 4K 1.22e+18 3.3994 29.95 BlockShuffle 3.0 6K 1.69e+18 3.3218 27.71 BlockShuffle 2.2 5K 1.69e+18 3.3011 27.14 Transformer-m 335.08 201.33 788.7 6.7 13K 1.55e+19 2.9062 18.29 LowRank 202.43 68.68 517.0 6.7 13K 1.01e+19 3.0251 20.60 LowRank 10.2 19K 1.54e+19 2.9359 18.84 LowRank 6.7 18K 1.55e+19 2.9310 18.75 BlockDense 198.67 64.91 509.3 6.7 13K 1.00e+19 3.0371 20.85 BlockDense 10.4 20K 1.55e+19 2.9491 19.09 BlockDense 6.7 18K 1.55e+19 2.9420 18.95 BlockShuffle 202.43 68.68 517.0 6.7 13K 1.01e+19 3.0501 21.12 BlockShuffle 10.2 19K 1.54e+19 2.9627 19.35 BlockShuffle 6.7 18K 1.55e+19 2.9525 19.15 Transformer-l 729.11 452.98 1646.9 14.6 28K 7.03e+19 2.6594 14.29 LowRank 430.66 154.53 1035.6 14.6 28K 4.42e+19 2.7527 15.69 LowRank 23.3 44K 7.03e+19 2.6917 14.76 LowRank 14.6 39K 7.01e+19 2.6850 14.66 BlockDense 430.66 154.53 1035.6 14.6 28K 4.42e+19 2.7570 15.75 BlockDense 23.3 44K 7.03e+19 2.6946 14.80 BlockDense 14.6 39K 7.01e+19 2.6941 14.79 BlockShuffle 430.66 154.53 1035.6 14.6 28K 4.42e+19 2.7735 16.01 BlockShuffle 23.3 44K 7.03e+19 2.7053 14.96 BlockShuffle 14.6 39K 7.01e+19 2.7104 15.04 Transformer-xl 1274.14 805.31 2814.3 25.5 49K 2.10e+20 2.5226 12.46 LowRank 743.56 274.73 1727.7 25.5 49K 1.29e+20 2.6062 13.55 LowRank 41.5 79K 2.10e+20 2.5464 12.76 LowRank 25.5 70K 2.10e+20 2.5539 12.86 BlockDense 728.49 259.65 1696.8 25.5 49K 1.27e+20 2.6204 13.74 BlockDense 42.2 80K 2.10e+20 2.5590 12.92 BlockDense 25.5 71K 2.10e+20 2.5637 12.98 BlockShuffle 743.56 274.73 1727.7 25.5 49K 1.29e+20 2.6254 13.81 BlockShuffle 41.5 79K 2.10e+20 2.5623 12.97 BlockShuffle 25.5 70K 2.10e+20 2.5678 13.03
(a) BlockDense
(b) BlockShuffle
Figure 12: Performance between dense Transformer, Structured FFN (BlockDense and BlockShuffle, LowRank in 1(b)) with 32% parameters either trained with self-guided training or more tokens across four sizes. Circle size indicates model FLOPs. To enable straightforward comparison, we controlled their training FLOPs to be the same.
Appendix C Implementation details
Model We employ the basic decoder-only Transformer architecture [41, 2] but enhance it with advanced components like Rotary Embedding [42] and SentencePiece Tokenizer [43], as used in recent Llama models [51]. We use a vocabulary size of 32,000 and enable tied word embeddings. The model employs layer normalization and the GELU activation function, with biases included in the architecture. We use random weight gaussian initialization with an initializer range of 0.02.
Structured FFN We limit the replacement of efficient linear layers only to FFN modules. This is because the linear layers in the attention module serve specific purposes, which can complicate analyses, and their efficient alternatives have been well-studied in the literature [52, 8, 22].
Dataset We use the RefinedWeb dataset [44], a carefully curated subset of CommonCrawl, processed for optimal filtering and deduplication. It provides 600B tokens available for public use. As it is a substantial dataset, we shuffle, extract, and tokenize it in advance. To manage memory efficiently, the token IDs are stored using np.memmap, preventing the need to simultaneously load all data into CPU memory. We set the maximum sequence length to be 1024. Obeying these, we randomly select 0.5B tokens for validation and reserve additional tokens for training. Following scaling laws [45], we allocate tokens at 20 times the number of each baseline model’s parameters.
Training We use A100 80G GPU to do the training and evaluation and adopt mixed precision technique (bfloat16 and float32) supported by torch.cuda.amp to accelerate training. For details, we employ the AdamW optimizer with 0.1 weight decay and [0.9, 0.999] betas and use the cosine annealing learning rate scheduler with 10% warm-up and $0.1$ times the minimum value. Learning rates are first swept for baseline models and applied consistently in subsequent experiments. Dropout is set to 0.0, and gradient clipping to 1.0. We measure the training FLOPs like Megatron [48], including all matrix multiplications. The configuration of each baseline model is put in Appendix C.
Table 11: Model and Training configuration. We report the number of layers (#Layer), hidden states dimension (Width), training tokens (Tokens)), global batch size in number of tokens (Batch), peaking learning rate (LR), and total training steps (Steps).
Name Size Width Layers Tokens Batch LR Steps Training FLOPs Transformer-s 110M 768 12 2.2B 0.5M 6.0e-4 4.2K 1.69e+18 Transformer-m 335M 1024 24 6.7B 0.5M 3.0e-4 13K 1.55e+19 Transformer-l 729M 1536 24 14.6B 0.5M 2.5e-4 28K 7.03e+19 Transformer-xl 1274M 2048 24 25.5B 0.5M 2.0e-4 49K 2.10e+20
Efficiency To boost efficiency in training and inference, our code is based on PyTorch but incorporates fast CUDA kernels. We utilize Flash Attention [24], fast LayerNorm, and rotary embedding from TransformerEngine [53], and fuse operations including bias and GeLU. For our inference speed test, we use bfloat16. These techniques are applied to all models to ensure a fair latency or throughput comparison.
Wide and structured network To reduce the parameters and FLOPs of both attention and FFN, we adopt a straightforward approach. First, the dimension of the Key or Value in attention is set to 256 following [7]. Second, the rank and attention dimension are selected to be half of the hidden state dimension. This creates a structured network that looks like a bottleneck. We put the detailed configuration in Table 12.
Table 12: The intermediate dim. refers to the intermediate dimension of FFN. Attention dim. refers to the dimension of scaled-dot product. KV dim indicates the dimension used for KVCache, which is picked according to Team et al. [7].
Method hidden dim. intermediate dim. attention dim. KV dim. Transformer-m 1024 4096 1024 256 Transformer-m (GQA) 1024 4864 1024 256 Ours (R=512) 1024 4864 512 256 Transformer-l 1536 6144 1536 256 Transformer-l (GQA) 1536 7424 1536 256 Ours (R=768) 1536 7424 768 256
Appendix D Broader Impacts
Enhancing the efficiency of Large Language Models (LLMs) can significantly reduce computational resources and energy consumption, benefiting the environment and democratizing access to advanced AI technologies. However, increased efficiency could also lead to greater dissemination of disinformation and the creation of deepfakes, posing risks to public trust and security and potentially reinforcing existing biases that impact specific groups unfairly. This research aims to promote the responsible development and deployment of LLMs, maximizing societal benefits while acknowledging potential harms.